OpenStudy (anonymous):

48= -16t^2+32t+48 solve for t

OpenStudy (anonymous):

t=0 and t=2

myininaya (myininaya):

subtract 48 on both sides 0=-16t^2+32t factor 0=-16t(t-2) -16t=0 so t=0 t-2=0 so t=2 gj

OpenStudy (anonymous):

thank you..